4. In the 2-dimensional space (Xo , xi), consider the coordinate transformation: (β is a parameter) Show that if A (40,A) & B (Bo, B1) are vectors with respect to the coordinate transformation above, then the scalar product: is invariant under the transformation (1) (i.e. it is indeed a scalar)
Answer #1

\boldsymbol{A'}\equiv (A_o',A_1') ; \boldsymbol{B'}\equiv (B_o',B_1') are the transformed vectors

A_o' = \gamma (A_o-\beta A_1) ; B_o' = \gamma (B_o-\beta B_1) \\ A_1' = \gamma (A_1-\beta A_o) ; B_o' = \gamma (B_1-\beta B_o)

A_o'B_o' = \gamma ^2\left [ A_oB_o+\beta ^2A_1B_1 - \beta (A_1B_o+A_oB_1) \right ]

A_1'B_1' = \gamma ^2\left [ A_1B_1+\beta ^2A_oB_ - \beta (A_1B_o+A_oB_1) \right ]

\boldsymbol{A'.B'} \equiv (A_o'B_o' - A_1'B_1')

  = \gamma ^2\left [ A_oB_o - A_1B_1- \beta ^2(A_oB_o - A_1B_1) \right ] \\ = \gamma ^2(1-\beta ^2)( A_oB_o - A_1B_1) \\ = ( A_oB_o - A_1B_1) \equiv \boldsymbol{A.B}

=> the scalar product is invariant under transformation